# From Models to Network Topologies: A Topology Inference Attack in Decentralized Federated Learning

### v2

@@ Abstract
Federated Learning (FL) is widely recognized as a privacy-preserving machine learning paradigm due to its model-sharing mechanism that avoids direct data exchange. Nevertheless, model training leaves exploitable traces that can be used to infer sensitive information. In Decentralized FL (DFL), the topology, defining how participants are connected, plays a crucial role in shaping the model’s privacy, robustness, and convergence. However, the topology introduces an unexplored vulnerability: attackers can exploit it to infer participant relationships and launch targeted attacks. This work uncovers the hidden risks of DFL topologies by proposing a novel Topology Inference Attack that infers the topology solely from model behavior. A taxonomy of topology inference attacks is introduced, categorizing them by the attacker’s capabilities and knowledge. Practical attack strategies are designed for various scenarios, and experiments are conducted to identify key factors influencing attack success. The results demonstrate that analyzing only the model of each node can accurately infer the DFL topology, highlighting a critical privacy risk in DFL systems. These findings offer valuable insights for improving privacy preservation in DFL environments.
@@ 1 Introduction
Therefore, this paper improves this research gap by proposing a novel overlay topology inference attack targeting DFL systems. The proposed attack leverages the behavioral traces generated by models to uncover sensitive overlay topologies. The main contributions of this work are as follows: (i) the formulation of a novel taxonomy of topology inference attacks, classifying these attacks based on the attacker’s capabilities and knowledge; (ii) the development of practical strategies tailored to various attack types, accompanied by a quantitative analysis of critical factors that influence the efficiency of these attacks; (iii) rigorous experimental evaluations conducted across a range of datasets and real-world network topology configurations to empirically validate the proposed attack strategies; and (iv) the provision of insights aimed at informing the design of effective defensive mechanisms to safeguard sensitive information and preserve privacy within DFL systems.
@@ 2.1 Inference Attacks
These attacks not only compromise users’ sensitive information and undermine data security but also erode trust in FL’s privacy protection mechanisms. As a result, users may hesitate to participate in the training process, undermining the overall effectiveness of the FL system. However, existing research on inference attacks in FL primarily focuses on CFL, with limited exploration of privacy leakage in DFL. Moreover, these studies have not adequately addressed the critical role of DFL’s overlay topology.
@@ 2.2.1 Communication Systems
In communication systems, most protocols are designed to avoid ring formation, resulting in tree-like overlay topologies. Moreover, many topology inference studies commonly assume a single path between nodes [5]. A widely used approach involves leveraging ICMP-based protocols, such as the traceroute, to map network topologies effectively [21]. Tomography-based metrics, like packet loss rates, further aid in deducing a system’s topology: by observing loss rates from a source node to a target node, one can estimate the likelihood of a direct connection and thereby reconstruct the network [6]. Additionally, by measuring the timestamps of sent and received packets, it is possible to determine the number of forwarding hops, offering deeper insight into the overall network configuration [19].
@@ 2.2.2 Social Networks
Social networks generally refer to the social structure formed by individuals in society connected by a specific relationship. In social networks, user interactions (e.g., likes, retweets) serve as valuable signals for topology modeling and inference by capturing how information diffuses across the network [13]. However, this behavioral data is often incomplete, especially in large-scale settings, necessitating methods like noisy sparse subspace clustering to infer network structures from partial observations [33]. Meanwhile, social networks are dynamic and time-sensitive, and they frequently exhibit small-world properties.[15] incorporate these distributional patterns as prior knowledge in evolutionary models to better capture network changes over time.
Communication networks that often assume tree topologies and social networks are typically modeled as small-world networks; however, these assumptions do not hold for DFL network topologies. Moreover, common inference tools and metrics, such as traceroute, packet loss rate, or user interaction data, are not readily applicable in DFL contexts, creating the need for specialized inference metrics and strategies tailored to the particularities of DFL systems.
@@ 3 Overlay Topology in DFL
where $\delta_{t}$ represents local updates, and $\tilde{M}_{t-1}$ is the aggregated model from round $t-1$.
@@
Given the critical role of overlay topology, it is essential to study potential information leakage through model updates and develop effective methods to protect sensitive topology information in DFL.
@@ 4 Problem Statement
For Scenario 1, since the attacker can control all nodes and knows partial information about the edges, it can be classified as a white-box attack. Scenarios 2, 3, and 4 can be categorized as gray-box attacks. For attack scenario 5, no information is available regarding the total number of nodes or edges, making it infeasible to fully reconstruct the network topology with the approaches considered. As a result, this work only considers the attack scenarios 1, 2, 3, and 4.
@@ 5.1 Attack Metrics
    Cosine $\text{Cosine Similarity}(\mathbf{a},\mathbf{b})=\frac{\mathbf{a}\cdot\mathbf{b}}{\|\mathbf{a}\|\,\|\mathbf{b}\|}$ (11)
@@
    Euclidean $\text{Euclidean Similarity}(\mathbf{a},\mathbf{b})=\frac{1}{1+\sqrt{\sum_{i=1}^{n}(a_{i}-b_{i})^{2}}}$ (12)
@@
    Curvature $\text{Curvature Divergence}(\Delta\mathbf{a},\Delta\mathbf{b})=\frac{\|\Delta\mathbf{a}-\Delta\mathbf{b}\|_{2}}{\frac{1}{2}(\|\Delta\mathbf{a}\|_{2}+\|\Delta\mathbf{b}\|_{2})}$ (13)
    where $\Delta\mathbf{a}=\mathbf{a}^{t}-\mathbf{a}^{t-1}$ and $\Delta\mathbf{b}=\mathbf{b}^{t}-\mathbf{b}^{t-1}$ denote parameter updates between rounds $t-1$ and $t$. A lower value indicates higher curvature divergence.
@@ 5.2 Attack Strategies
Given $N$ nodes and their behavioral metrics $X\in\mathbb{R}^{N\times d}$, EDGEPRE directly uses an MLP-based edge classifier without any message passing. For each labeled pair $(i,j)$, the concatenated feature pair $[x_{i}\|x_{j}]$ is passed to a multilayer perceptron (MLP) decoder that outputs the predicted edge probability:
@@
5:     $Z\leftarrow f_{\text{enc}}(X)$triangleright Node embeddings via multi-head GAT
6:     $A^{\prime}_{ij}\leftarrow f_{\text{dec}}(z_{i},z_{j}),\quad\forall i,j$triangleright Decoder predicts edge weights
7:     $A^{\prime}\leftarrow\frac{1}{2}(A^{\prime}+A^{\prime\top})$triangleright Symmetrize adjacency
8:     $\mathcal{L}\leftarrow\sum_{i,j}\left\|A^{\prime}_{ij}-X_{ij}\right\|^{2}$triangleright Reconstruction loss
9:     $\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}\mathcal{L}$triangleright Gradient descent update
@@ 6 Evaluation
This work empirically validates the effectiveness of the proposed topology inference attack. First, it identifies which metrics yield stronger attack performance. Building on these results, it then thoroughly evaluates the proposed attack strategies under various conditions, including different datasets, topologies, node counts, and data preprocessing methods. This comprehensive assessment demonstrates the robustness and applicability of the proposed topology inference attack.
@@ 6.1 Experimental Setups
A. Dataset and Model: All experiments are conducted on five datasets: MNIST, Fashion-MNIST, CIFAR10, SVHM, and ImageNet100, widely used in FL and inference attack benchmarks.
  • MNIST [23]: Contains 60,000 grayscale training images and 10,000 test images of handwritten digits (0–9) at 28$\times$28 resolution. A two-hidden-layer MLP (256 and 128 neurons) trained with Adam (lr=1e-3) is used.
  • FMNIST [35]: Similar size and format to MNIST, but with 10 categories of clothing items. A CNN with two convolutional layers (32 and 64 filters, kernel size 3$\times$3) is employed.
  • CIFAR10 [22]: Contains 60,000 32$\times$32 RGB images from 10 categories. The training set is only treated with basic image normalization. The image pixel values are normalized to keep the original image content unchanged. A MobileNet [30] is used for training.
  • SVHN [26]: Composes of over 600,000 32$\times$32 RGB images of house numbers collected from Google Street View. Each image contains a centered digit from 0 to 9. A ResNet9 [17] is adopted as the training model.
@@
B. DFL Topology Setting: All experiments are conducted on DFL systems comprising two categories of topologies: synthetic and real-world. The synthetic topologies include ring, star, and three Erdős-Rényi (ER) graphs with connection probabilities $p=$0.3, 0.5, and 0.7 [9]. The real-world topologies consist of 29 network topologies collected from the SNDlib [27] and DEFO [16] repositories, which reflect realistic communication and infrastructure networks. Detailed descriptions and properties of these topologies are provided in the supplementary material.
@@ C. Federation Setting: All experiments employ the following DFL configurations:
  • Number of Nodes: For synthetic topologies, the number of nodes is set to 10, 20, and 30. For real-world topologies, the number of nodes ranges from 12 to 161, depending on the specific topology.
@@
D. Attack Models: To evaluate the effectiveness of the proposed methods, EDGEPRE is compared against classical supervised classifiers, including Logistic Regression [34], Support Vector Machine (SVM) [18], and Random Forest (RF) [4]. For the unsupervised attack strategy, INFERGAT is evaluated against standard clustering-based methods, including K-Means [24], Gaussian Mixture Model (GMM) [28], and Spectral Clustering [32].
E. Evaluation Metric: Since the topology inference attack is formulated as a binary classification problem, the F1-Score set is used as the evaluation metric.

### v3

@@ Abstract
Federated Learning (FL) is widely recognized as a privacy-preserving Machine Learning paradigm due to its model-sharing mechanism that avoids direct data exchange. Nevertheless, model training leaves exploitable traces that can be used to infer sensitive information. In Decentralized FL (DFL), the topology, defining how participants are connected, plays a crucial role in shaping the model’s privacy, robustness, and convergence. However, the topology introduces an unexplored vulnerability: attackers can exploit it to infer participant relationships and launch targeted attacks. This work uncovers the hidden risks of DFL topologies by proposing a novel Topology Inference Attack that infers the topology solely from model behavior. A taxonomy of topology inference attacks is introduced, categorizing them by the attacker’s capabilities and knowledge. Practical attack strategies are designed for various scenarios, and experiments are conducted to identify key factors influencing attack success. The results demonstrate that analyzing only the model of each node can accurately infer the DFL topology, highlighting a critical privacy risk in DFL systems. These findings offer insights for improving privacy preservation in DFL environments.
@@ 1 Introduction
Therefore, this paper improves this research gap by proposing a novel overlay topology inference attack targeting DFL systems. The proposed attack leverages the behavioral traces generated by models to uncover sensitive overlay topologies. The main contributions of this work are as follows: (i) the formulation of a novel taxonomy of topology inference attacks, classifying these attacks based on the attacker’s capabilities and knowledge; (ii) the development of practical strategies tailored to various attack types, accompanied by a quantitative analysis of critical factors that influence the efficiency of these attacks  [^1] Code available at: https://github.com/luke-feng/TIA; (iii) rigorous experimental evaluations conducted across a range of datasets and real-world network topology configurations to empirically validate the proposed attack strategies; and (iv) the provision of insights aimed at informing the design of effective defensive mechanisms to safeguard sensitive information within DFL systems.
@@ 2.1 Inference Attacks
These attacks not only compromise users’ sensitive information and undermine data security but also erode trust in FL’s privacy preserving mechanisms. Users may hesitate to participate in the training process, undermining the overall effectiveness of the FL system. However, existing research primarily focuses on CFL, with limited exploration of privacy leakage in DFL. Moreover, these studies have not adequately addressed the critical role of DFL’s overlay topology.
@@ 2.2.1 Communication Systems
In communication systems, most protocols are designed to avoid ring formation, resulting in tree-like topologies. Many topology inference studies commonly assume a single path between nodes [3]. A widely used approach involves leveraging ICMP-based protocols, such as the traceroute, to map network topologies [19]. Tomography-based metrics, like packet loss rates, further aid in deducing a system’s topology: by observing loss rates from a source node to a target node, one can estimate the likelihood of a direct connection and thereby reconstruct the network [4]. Additionally, by measuring the timestamps of sent and received packets, it is possible to determine the number of forwarding hops, offering deeper insight into the overall network configuration [17].
@@ 2.2.2 Social Networks
Social networks generally refer to the social structure formed by individuals in society connected by a specific relationship. In social networks, user interactions (e.g., likes, retweets) serve as signals for topology modeling and inference by capturing how information diffuses across the network [11]. However, this behavioral data is often incomplete, especially in large-scale settings, necessitating methods like noisy sparse subspace clustering to infer network structures from partial observations [30]. Meanwhile, social networks are dynamic and time-sensitive, and they frequently exhibit small-world properties.[13] incorporate these distributional patterns as prior knowledge in evolutionary models to better capture network changes over time.
To conclude, communication networks often assume tree topologies and social networks are typically modeled as small-world networks. These assumptions do not hold for DFL network topologies. Moreover, common inference tools and metrics, such as traceroute, packet loss rate, or user interaction data, are not readily applicable in DFL contexts, creating the need for specialized inference metrics and strategies tailored to the particularities of DFL systems.
@@ 3 Overlay Topology in DFL
where $\delta_{t}$ represents local updates, and $\tilde{M}_{t-1}$ is the aggregated model from round t1t-1.
@@
Given the critical role of overlay topology, it is essential to investigate potential information leakage through model updates and develop effective methods to protect sensitive topology information in DFL.
@@ 4 Problem Statement
For Scenario 1, since the attacker can control all nodes and knows partial information about the edges, it can be classified as a white-box attack. Scenarios 2, 3, and 4 can be categorized as gray-box attacks. For attack scenario 5, no information is available regarding the total number of nodes or edges, making it difficult to fully reconstruct the network topology.
@@ 5.1 Attack Metrics
    $\text{Cosine Similarity}(\mathbf{a},\mathbf{b})=\frac{\mathbf{a}\cdot\mathbf{b}}{\|\mathbf{a}\|\,\|\mathbf{b}\|}$ (11)
@@
    $\text{Euclidean Similarity}(\mathbf{a},\mathbf{b})=\frac{1}{1+\sqrt{\sum_{i=1}^{n}(a_{i}-b_{i})^{2}}}$ (12)
@@
    $\text{Curvature Divergence}(\Delta\mathbf{a},\Delta\mathbf{b})=\frac{\|\Delta\mathbf{a}-\Delta\mathbf{b}\|_{2}}{\frac{1}{2}(\|\Delta\mathbf{a}\|_{2}+\|\Delta\mathbf{b}\|_{2})}$ (13)
    where $\Delta\mathbf{a}=\mathbf{a}^{t}-\mathbf{a}^{t-1}$ and $\Delta\mathbf{b}=\mathbf{b}^{t}-\mathbf{b}^{t-1}$ denote parameter updates between rounds t1t-1 and $t$. A lower value indicates lower divergence (i.e., more similar updates).
@@ 5.2 Attack Strategies
Given $N$ nodes and their behavioral metrics $X\in\mathbb{R}^{N\times d}$, EDGEPRE directly uses an MLP-based edge classifier without any message passing. For each labeled pair $(i,j)$, the concatenated feature pair $x_{i}\|x_{j}]$ is passed to a multilayer perceptron (MLP) decoder that outputs the predicted edge probability:
@@
5:  $Z\leftarrow f_{\text{enc}}(X)$ $\triangleright$ Node embeddings via multi-head GAT
6:  $A^{\prime}_{ij}\leftarrow f_{\text{dec}}(z_{i},z_{j}),\quad\forall i,j$ $\triangleright$ Decoder predicts edge weights
7:  $A^{\prime}\leftarrow\frac{1}{2}(A^{\prime}+A^{\prime\top})$ $\triangleright$ Symmetrize adjacency
8:  $\mathcal{L}\leftarrow\sum_{i,j}\left\|A^{\prime}_{ij}-X_{ij}\right\|^{2}$ $\triangleright$ Reconstruction loss
9:  $\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}\mathcal{L}$ $\triangleright$ Gradient descent update
@@ 6 Evaluation
This work empirically validates the effectiveness of the proposed topology inference attack. First, it identifies which metrics yield stronger attack performance. Building on these results, it then thoroughly evaluates the proposed attack strategies under various conditions, including different datasets, topologies, node counts, and data preprocessing methods.
@@ 6.1 Experimental Setups
A. Dataset and Model: Experiments are conducted on: MNIST, FMNIST, CIFAR10, SVHN, ImageNet10, and Malware, which are widely used in FL and inference attack benchmarks.
  • MNIST [21]: Contains 60,000 grayscale training images and 10,000 test images of handwritten digits (0–9) at $\times28$ resolution. A two-hidden-layer MLP (256 and 128 neurons) trained with Adam (lr=1e-3) is used.
  • FMNIST [31]: Similar size and format to MNIST, but with 10 categories of clothing items. A CNN with two convolutional layers (32 and 64 filters, kernel size $\times3$) is employed.
  • CIFAR10 [20]: Contains 60,000 $\times32$ RGB images from 10 categories. The training set is only treated with basic image normalization. The image pixel values are normalized to keep the original image content unchanged. A MobileNet [26] is used for training.
  • SVHN [23]: Composes of over 600,000 $\times32$ RGB images of house numbers collected from Google Street View. Each image contains a centered digit from 0 to 9. A ResNet9 [16] is adopted as the training model.
@@
  • Malware [8]: The Malware dataset comprises 342,106 tabular records with 31 extracted features, collected from eight IoT crowdsensing devices infected by eight types of malware. Each record is labeled as either the corresponding malware type or normal behavior. A MLP with two hidden layers (64 and 32 neurons) is employed.
B. DFL Topology Setting: All experiments are conducted on DFL systems comprising two categories of topologies: synthetic and real-world. The synthetic topologies include ring, star, and three Erdős-Rényi (ER) graphs with connection probabilities p=p=0.3, 0.5, and 0.7. The real-world topologies consist of 27 network topologies collected from the SNDlib [24] and DEFO [15] repositories, which reflect realistic communication and infrastructure networks. Detailed descriptions and properties of these topologies are provided in the supplementary material.
@@ C. Federation Setting: All experiments employ the following DFL configurations:
  • Number of Nodes: For synthetic topologies, the number of nodes is set to 10, 20, and 30. For real-world topologies, the number of nodes ranges from 12 to 87, depending on the specific topology.
@@
D. Attack Models: To evaluate the effectiveness of the proposed methods, EDGEPRE is compared against classical supervised classifiers, including Logistic Regression, Support Vector Machine (SVM), and Random Forest (RF). Additionally, graph-specific supervised baselines are also used, including: SEAL [34] and GraphSAGE (edge classification setting) [14].
For the unsupervised attack strategy, INFERGAT is evaluated against standard clustering-based methods, including K-Means, Gaussian Mixture Model (GMM), and Spectral Clustering. This experiment further adds modern graph-reconstruction and self-supervised baselines: GRACE [35] and BGRL [29].
E. Evaluation Metric: Since the topology inference attack is formulated as a binary classification problem, evaluation uses the F1-Score and AUC-ROC metrics.
@@ 6.2 Selection of Attack Metrics
Scenario Algorithm CIFAR10 FMNIST ImageNet10 Malware MNIST SVHN
F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC
SC1 Logistic 0.806 $\pm$ 0.129 0.519 $\pm$ 0.064 0.804 $\pm$ 0.126 0.500 $\pm$ 0.000 0.804 $\pm$ 0.126 0.500 $\pm$ 0.000 0.804 $\pm$ 0.126 0.500 $\pm$ 0.000 0.800 $\pm$ 0.125 0.500 $\pm$ 0.000 0.804 $\pm$ 0.126 0.500 $\pm$ 0.000
RF 0.795 $\pm$ 0.132 0.643 $\pm$ 0.127 0.786 $\pm$ 0.142 0.653 $\pm$ 0.128 0.833 $\pm$ 0.107 0.695 $\pm$ 0.132 0.714 $\pm$ 0.144 0.503 $\pm$ 0.050 0.765 $\pm$ 0.135 0.600 $\pm$ 0.102 0.788 $\pm$ 0.114 0.633 $\pm$ 0.116
SVM 0.840 $\pm$ 0.102 0.609 $\pm$ 0.153 0.818 $\pm$ 0.115 0.561 $\pm$ 0.122 0.808 $\pm$ 0.118 0.519 $\pm$ 0.066 0.800 $\pm$ 0.137 0.499 $\pm$ 0.025 0.801 $\pm$ 0.131 0.519 $\pm$ 0.067 0.824 $\pm$ 0.108 0.566 $\pm$ 0.130
GraphSage 0.740 $\pm$ 0.077 0.607 $\pm$ 0.253 0.699 $\pm$ 0.058 0.474 $\pm$ 0.188 0.701 $\pm$ 0.045 0.507 $\pm$ 0.209 0.686 $\pm$ 0.043 0.423 $\pm$ 0.186 0.686 $\pm$ 0.039 0.460 $\pm$ 0.185 0.712 $\pm$ 0.056 0.526 $\pm$ 0.200
SEAL 0.701 $\pm$ 0.057 0.560 $\pm$ 0.142 0.697 $\pm$ 0.052 0.556 $\pm$ 0.152 0.699 $\pm$ 0.057 0.557 $\pm$ 0.156 0.705 $\pm$ 0.052 0.577 $\pm$ 0.144 0.701 $\pm$ 0.050 0.584 $\pm$ 0.129 0.697 $\pm$ 0.056 0.548 $\pm$ 0.146
EDGEPRE 0.847 $\pm$ 0.101 0.831 $\pm$ 0.136 0.873 $\pm$ 0.096 0.857 $\pm$ 0.135 0.855 $\pm$ 0.101 0.826 $\pm$ 0.181 0.858 $\pm$ 0.102 0.834 $\pm$ 0.192 0.862 $\pm$ 0.095 0.839 $\pm$ 0.144 0.886 $\pm$ 0.091 0.871 $\pm$ 0.130
SC2 Logistic 0.811 $\pm$ 0.128 0.543 $\pm$ 0.101 0.804 $\pm$ 0.126 0.500 $\pm$ 0.000 0.801 $\pm$ 0.127 0.500 $\pm$ 0.001 0.804 $\pm$ 0.126 0.500 $\pm$ 0.000 0.800 $\pm$ 0.125 0.500 $\pm$ 0.000 0.804 $\pm$ 0.126 0.500 $\pm$ 0.000
RF 0.787 $\pm$ 0.134 0.629 $\pm$ 0.132 0.762 $\pm$ 0.135 0.568 $\pm$ 0.107 0.784 $\pm$ 0.122 0.579 $\pm$ 0.096 0.722 $\pm$ 0.139 0.505 $\pm$ 0.048 0.746 $\pm$ 0.140 0.562 $\pm$ 0.082 0.765 $\pm$ 0.129 0.598 $\pm$ 0.102
SVM 0.836 $\pm$ 0.108 0.604 $\pm$ 0.151 0.817 $\pm$ 0.117 0.559 $\pm$ 0.119 0.810 $\pm$ 0.116 0.520 $\pm$ 0.067 0.800 $\pm$ 0.137 0.499 $\pm$ 0.025 0.800 $\pm$ 0.132 0.517 $\pm$ 0.066 0.821 $\pm$ 0.110 0.560 $\pm$ 0.128
GraphSage 0.697 $\pm$ 0.050 0.488 $\pm$ 0.172 0.700 $\pm$ 0.069 0.510 $\pm$ 0.152 0.699 $\pm$ 0.050 0.501 $\pm$ 0.135 0.700 $\pm$ 0.056 0.473 $\pm$ 0.157 0.698 $\pm$ 0.056 0.498 $\pm$ 0.139 0.699 $\pm$ 0.060 0.490 $\pm$ 0.152
SEAL 0.699 $\pm$ 0.042 0.558 $\pm$ 0.137 0.704 $\pm$ 0.052 0.582 $\pm$ 0.108 0.700 $\pm$ 0.046 0.589 $\pm$ 0.105 0.700 $\pm$ 0.050 0.561 $\pm$ 0.145 0.701 $\pm$ 0.049 0.570 $\pm$ 0.133 0.706 $\pm$ 0.051 0.579 $\pm$ 0.130
EDGEPRE 0.786 $\pm$ 0.093 0.721 $\pm$ 0.175 0.825 $\pm$ 0.108 0.791 $\pm$ 0.168 0.811 $\pm$ 0.091 0.772 $\pm$ 0.153 0.822 $\pm$ 0.112 0.778 $\pm$ 0.188 0.827 $\pm$ 0.108 0.797 $\pm$ 0.161 0.826 $\pm$ 0.109 0.784 $\pm$ 0.167
SC3 BGRL 0.384 $\pm$ 0.344 0.500 $\pm$ 0.000 0.384 $\pm$ 0.344 0.500 $\pm$ 0.000 0.384 $\pm$ 0.344 0.500 $\pm$ 0.000 0.384 $\pm$ 0.344 0.500 $\pm$ 0.000 0.433 $\pm$ 0.349 0.500 $\pm$ 0.000 0.384 $\pm$ 0.344 0.500 $\pm$ 0.000
GMM 0.743 $\pm$ 0.252 0.674 $\pm$ 0.081 0.742 $\pm$ 0.201 0.783 $\pm$ 0.129 0.668 $\pm$ 0.227 0.731 $\pm$ 0.161 0.795 $\pm$ 0.107 0.748 $\pm$ 0.136 0.864 $\pm$ 0.150 0.890 $\pm$ 0.156 0.852 $\pm$ 0.186 0.869 $\pm$ 0.126
GRACE 0.866 $\pm$ 0.000 0.426 $\pm$ 0.207 0.534 $\pm$ 0.469 0.446 $\pm$ 0.181 0.866 $\pm$ 0.000 0.396 $\pm$ 0.207 0.500 $\pm$ 0.337 0.441 $\pm$ 0.189 0.534 $\pm$ 0.469 0.438 $\pm$ 0.186 0.866 $\pm$ 0.000 0.429 $\pm$ 0.210
Kmeans 0.793 $\pm$ 0.090 0.636 $\pm$ 0.087 0.741 $\pm$ 0.187 0.891 $\pm$ 0.134 0.792 $\pm$ 0.223 0.774 $\pm$ 0.181 0.779 $\pm$ 0.101 0.764 $\pm$ 0.164 0.754 $\pm$ 0.178 0.820 $\pm$ 0.152 0.907 $\pm$ 0.109 0.934 $\pm$ 0.093
Spectral 0.727 $\pm$ 0.168 0.592 $\pm$ 0.137 0.710 $\pm$ 0.268 0.594 $\pm$ 0.189 0.684 $\pm$ 0.314 0.709 $\pm$ 0.172 0.875 $\pm$ 0.111 0.721 $\pm$ 0.207 0.677 $\pm$ 0.231 0.594 $\pm$ 0.189 0.760 $\pm$ 0.148 0.569 $\pm$ 0.146
INFERGAT 0.893 $\pm$ 0.103 0.876 $\pm$ 0.161 0.850 $\pm$ 0.116 0.789 $\pm$ 0.147 0.765 $\pm$ 0.158 0.813 $\pm$ 0.157 0.664 $\pm$ 0.154 0.793 $\pm$ 0.158 0.657 $\pm$ 0.199 0.860 $\pm$ 0.186 0.769 $\pm$ 0.184 0.692 $\pm$ 0.391
SC4 BGRL 0.368 $\pm$ 0.243 0.500 $\pm$ 0.000 0.368 $\pm$ 0.243 0.500 $\pm$ 0.000 0.368 $\pm$ 0.243 0.500 $\pm$ 0.000 0.368 $\pm$ 0.243 0.500 $\pm$ 0.000 0.374 $\pm$ 0.242 0.500 $\pm$ 0.000 0.368 $\pm$ 0.243 0.500 $\pm$ 0.000
GMM 0.360 $\pm$ 0.116 0.292 $\pm$ 0.129 0.328 $\pm$ 0.137 0.274 $\pm$ 0.149 0.335 $\pm$ 0.163 0.324 $\pm$ 0.160 0.331 $\pm$ 0.141 0.309 $\pm$ 0.146 0.303 $\pm$ 0.150 0.270 $\pm$ 0.157 0.299 $\pm$ 0.133 0.255 $\pm$ 0.149
GRACE 0.394 $\pm$ 0.244 0.536 $\pm$ 0.191 0.394 $\pm$ 0.244 0.522 $\pm$ 0.188 0.393 $\pm$ 0.241 0.505 $\pm$ 0.190 0.392 $\pm$ 0.244 0.510 $\pm$ 0.194 0.401 $\pm$ 0.247 0.538 $\pm$ 0.190 0.395 $\pm$ 0.248 0.548 $\pm$ 0.187
Kmeans 0.389 $\pm$ 0.097 0.272 $\pm$ 0.122 0.373 $\pm$ 0.125 0.234 $\pm$ 0.129 0.359 $\pm$ 0.136 0.280 $\pm$ 0.155 0.371 $\pm$ 0.121 0.269 $\pm$ 0.135 0.356 $\pm$ 0.115 0.240 $\pm$ 0.144 0.344 $\pm$ 0.128 0.230 $\pm$ 0.141
Spectral 0.448 $\pm$ 0.136 0.448 $\pm$ 0.136 0.439 $\pm$ 0.110 0.439 $\pm$ 0.110 0.465 $\pm$ 0.141 0.465 $\pm$ 0.141 0.443 $\pm$ 0.118 0.443 $\pm$ 0.118 0.456 $\pm$ 0.116 0.456 $\pm$ 0.116 0.435 $\pm$ 0.169 0.435 $\pm$ 0.169
INFERGAT 0.643 $\pm$ 0.194 0.790 $\pm$ 0.140 0.604 $\pm$ 0.232 0.781 $\pm$ 0.132 0.594 $\pm$ 0.201 0.771 $\pm$ 0.134 0.607 $\pm$ 0.179 0.757 $\pm$ 0.148 0.660 $\pm$ 0.185 0.794 $\pm$ 0.145 0.673 $\pm$ 0.199 0.805 $\pm$ 0.152
Table 1: F1 and AUC (mean $\pm$ std) across datasets. Best per scenario/dataset in bold.